OpenStudy (flowey15):

A tree that is 100 feet tall casts a shadow that is 150 feet long. Determine the angle at which the rays of the sun hit the ground, to the nearest degree. A) 31° B) 34° C) 42° D) 56°

OpenStudy (math_man21):

is this pythagorean theorem?

OpenStudy (flowey15):

no,, because the answer isn't a choice

OpenStudy (error1603):

Welcome! \(\Huge\cal\rlap{\color{blue}{ERROR^lsa^21^3 \star}}{\color{lime}{\:ERROR^lsa^21^3 \star }}\)

OpenStudy (calculusxy):



OpenStudy (calculusxy):

We need to find what \(\theta\) would equal to. We do have the measurement of the side that is opposite and adjacent to the angle \(\theta\). That means that we need use \(\tan^{-1}\) because using the inverse tangent operation will help us find the angle measure of \(\theta\). So what you would do is \(\tan^{-1}(\frac{100}{150})\).

OpenStudy (flowey15):

i don't have my calculator, that's my problem

OpenStudy (calculusxy):

Okay. So you could google "calculator" and they would provide you with a calculator.

OpenStudy (calculusxy):

https://www.google.com/?gws_rd=ssl#q=calculator

OpenStudy (calculusxy):

Did you get it?

OpenStudy (flowey15):

no, i got this...........1.10714871779

OpenStudy (calculusxy):

So you should get 34 degrees.